1) the section shown is subject to a 300 N-m moment at an angle of β = 15o

a) What is the angle that the neutral axis lays on?
b) What is the maximum compression and tension?

c) what are the stresses at points A and B?

nt at ays M = 300 N·m 20 mm 10 mm 20 mm 10 mm
Answer #1

If you like the solution please give thumbs up

Soph B=15° M=300 N-m AMsing 20mm Mze=MCopß = 300 (oxigo M2=289.78 N-m 2 44 mlost a tomm My = Msinß 20mm - 300 sinis My=77.646 For maym Compression ма. at hy 5 + 5 + her as ty Iz = 77.64 x 183 X5 2.5834104 289.78x103x25 To 58x10 Marion Comp. [ 5A = -